OpenStudy (anonymous):

What is the value of the expression 3 – a – b when a = 15 and b = –3? A. –21 B. –15 C. –9 D. 21

OpenStudy (anonymous):

wat do u think

OpenStudy (anonymous):

let see... A?

OpenStudy (anonymous):

its c wanna know y

OpenStudy (welshfella):

3 - 15 - (-3)

OpenStudy (anonymous):

Why

OpenStudy (youngstudier):

@YK.PAPI , please stop giving out direct answers.

OpenStudy (anonymous):

if u do the math correctly fill in the blanks it equals c

OpenStudy (anonymous):

omg i asked wat she thought first jesus

OpenStudy (youngstudier):

*Sigh* 3 -15 = -12 + 3 = -9.

OpenStudy (anonymous):

XD

OpenStudy (youngstudier):

You're supposed to help them get the answer, not give it to them after asking what they thought it was.

OpenStudy (anonymous):

do u understand it and dis is to easy to work out so

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

I PASSED!!!

OpenStudy (anonymous):

good im glad you did

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

wat grade are u in

OpenStudy (anonymous):

7th wbu

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

Nice Just One More Year Of School Lol

OpenStudy (anonymous):

I got 5 to go

OpenStudy (anonymous):

lol ikr i cant wait and do u like skool

OpenStudy (anonymous):

Lol I hate School

OpenStudy (anonymous):

lol omg me to

OpenStudy (anonymous):

id think anyone like school Lol

OpenStudy (anonymous):

lol ik some ppl that love skool

OpenStudy (anonymous):

O_o Are you sure your from earth?

OpenStudy (anonymous):

im absolutely sure i am👽

OpenStudy (anonymous):

Lol Well i gtg bi

OpenStudy (anonymous):

bye lol

OpenStudy (anonymous):

Bye :3